# Hashed Watermark as a Filter: Defeating Forging and Overwriting Attacks in Weight-based Neural Network Watermarking

### v1

@@ VI-C Robustness Evaluation
Question 3. Is NeuralMark robust against overwriting attacks targeting watermark-specific layers, especially under varying attack strength levels?
We follow the setting outlined in Section IV-A to assess the robustness of NeuralMark against overwriting attacks. Specifically, we perform overwriting attacks targeting watermark-specific layers, with the number of training epochs set to 100 to simulate limited computational resources, and analyze the impact of two key factors: the hyper-parameter $\lambda$ in Eq. (2) and the learning rate $\eta$. Here, $\lambda$ controls the strength of the watermark embedding, with larger values leading to stronger embedding, while $\eta$ primarily affects model performance.
Distinct Values of $\lambda$. First, we investigate the influence of $\lambda$ in overwriting attacks. Specifically, we set $\lambda$ to 1, 10, 50, 100, and 1000, respectively. Table V presents the results on the CIFAR-100 to CIFAR-10 and CIFAR-10 to CIFAR-100 tasks using ResNet-18. We report only the original watermark detection rate, as the adversary’s watermark detection rate reaches $100\%$. As defined in the success criterion in Section IV-B, the original watermark must be effectively removed for overwriting attacks to be deemed successful. Thus, the overwriting attack experiments focus solely on whether the original watermark can be successfully removed. We can summarize several insightful observations. (1) As $\lambda$ increases, the original watermark detection rate of NeuralMark remains at 100%, while those of VanillaMark, GreedyMark, and VoteMark significantly decline. In particular, when $\lambda=1000$, the embedding strength of the adversary’s watermark is 1000 times greater than that of the original watermark. At this point, the original watermark detection rates for NeuralMark, VanillaMark, GreedyMark, and VoteMark on the CIFAR-100 to CIFAR-10 tasks are $100\%$, $53.90\%$, $49.60\%$, and 59.37%, respectively. Those results indicate that NeuralMark exhibits strong robustness against overwriting attacks. (2) As $\lambda$ increases, model performance remains relatively stable. This is because overwriting attacks jointly train both the main task and the watermark embedding task, enabling the model parameters to effectively adapt to both.
Distinct Values of $\eta$. Second, we examine the impact of $\eta$ in overwriting attacks. Concretely, we set $\eta$ to 0.001, 0.005, 0.01, 0.1, and 1, respectively. Table V lists the results on the CIFAR-100 to CIFAR-10 and CIFAR-10 to CIFAR-100 tasks using ResNet-18. We have the following important observations. (1) As $\eta$ increases, model performance declines due to its substantial impact on performance. Thus, the adversary cannot arbitrarily increase $\eta$ to strengthen the attack. (2) At $\eta=0.005$, the original watermark detection rates for VanillaMark, GreedyMark, and VoteMark drop dramatically, whereas NeuralMark maintains a detection rate close to $100\%$. (3) When $\eta=0.01$, the model performance of NeuralMark on the CIFAR-100 to CIFAR-10 task decreases by $2.07\%$, but its original watermark detection rate remains above the security boundary of 88.28% defined in Section V-C, while those for the other methods fall significantly. (4) For $\eta>=0.1$, although the original watermark detection rate of NeuralMark drops below the security boundary, the model performance is completely compromised, indicating that the attack is ineffective.
TABLE VI: Comparison of resistance to fine-tuning attacks. Values (%) inside and outside the bracket are the watermark detection rate and classification accuracy, respectively.
@@
TABLE VII: Comparison of resistance to fine-tuning attacks against watermark embedding layer using ResNet-18. Values (%) inside and outside the bracket are the watermark detection rate and classification accuracy, respectively.
Fine-tuning Clean NeuralMark VanillaMark GreedyMark VoteMark
@@
Figure 4: Comparison of resistance to pruning attacks under various pruning ratios on CIFAR-10 using AlexNet and ResNet-18: (a) NeuralMark, (b) VanillaMark, (c) GreedyMark, and (d) VoteMark.
Question 4. Is NeuralMark robust to fine-tuning attacks?
We follow the setting stated in Section IV-A to evaluate the robustness of NeuralMark against fine-tuning attacks, with training limited to 100 epochs to reflect realistic computational constraints.
Fine-tuning All Model Parameters. Following [14], we adopt the same hyper-parameters for fine-tuning attacks as during training, except for setting the learning rate to 0.001. Also, we replace the task-specific classifier with randomly initialized parameters and optimize all parameters by only minimizing the main task loss $\mathcal{L}_{m}$. Table VI reports the results of fine-tuning attacks, we find that watermarks embedded with NeuralMark maintain a $100\%$ watermark detection rate across all fine-tuning tasks. In contrast, watermarks embedded with VanillaMark, GreedyMark, and VoteMark experience a slight reduction in detection rates across several tasks. Those results indicate that fine-tuning attacks cannot effectively remove watermarks embedded with NeuralMark.
Fine-tuning Watermark-Specific Layer Parameters. Furthermore, we follow the aforementioned settings to perform fine-tuning attacks on the watermark-specific layers and the classifier, while keeping all other parameters frozen. As shown in Table VII, we can observe that the watermark detection rate remains at 100%, but the model performance of all methods exhibits a substantial decline. Specifically, for the CIFAR-10 to CIFAR-100 task using ResNet-18, the accuracy achieved by NeuralMark is 49.77%, which is markedly lower than the 71.67% accuracy obtained when all parameters are fine-tuned. Similar trends are observed across other methods. Those results indicate that fine-tuning only the watermark embedding layer and classifier makes it challenging to maintain effective model performance, failing fine-tuning attacks.
Question 5. Is NeuralMark robust to pruning attacks?
We now verify the robustness of NeuralMark in resisting pruning attacks. Specifically, we randomly reset a specified proportion of model parameters in the watermark embedding layer to zero. Fig. 4 shows the results of pruning attacks on the CIFAR-10 dataset using AlexNet and ResNet, respectively. As can be seen, as the pruning ratio increases, the performance of NeuralMark degrades while the detection rate remains nearly 100%. This indicates NeuralMark’s robustness against pruning attacks. Moreover, we observe that both VanillaMark and VoteMark exhibit strong resistance to pruning attacks, while GreedyMark demonstrates relatively weak resistance. One possible reason is that GreedyMark depends on several important parameters, and their removal may affect its robustness. More results of pruning attacks across distinct datasets are provided in Appendix B. All the results suggest that NeuralMark effectively resists pruning attacks.
VI-D Hashed Watermark Filter Analysis
Question 6. How does the number of filtering rounds impact the fidelity of NeuralMark?
To evaluate the impact of filtering rounds on NeuralMark’s fidelity, we conduct experiments with 6 and 8 filters, compared to the default 4 filters. Table VIII presents the impact of watermark embedding on the model performance across distinct filtering rounds. The results demonstrate that NeuralMark, even with varying filtering rounds, has a minimal effect on the model performance while successfully embedding watermarks.
@@ VI-D Hashed Watermark Filter Analysis
TABLE VIII: Comparison of classification accuracy (%) with various distinct filter rounds on CIFAR-10 and CIFAR-100 using ResNet-18. Watermark detection rates are omitted as they all reach 100%.
@@
Question 7. How does the number of filtering rounds influence the robustness of NeuralMark?
To assess the influence of the number of filtering rounds on NeuralMark’s robustness, we conduct experiments with 6 and 8 filters, compared to the default 4 filters. We omit forging attacks, as the hashed watermark filter is resilient to them.
TABLE IX: Comparison of resistance to overwriting attacks at various trade-off hyper-parameters ($\lambda$) and learning rates ($\eta$) with distinct filtering rounds using ResNet-18. Values (%) inside and outside the bracket are watermark detection rate and classification accuracy, respectively.
@@
Overwriting Attacks. Table IX lists the results of overwriting attacks across distinct filtering rounds. From the results, we find that when the number of filtering rounds is set to 6, NeuralMark exhibits superior robustness compared to 4 and 8 filter rounds. Specifically, at $\eta=0.01$, the original watermark detection rates for 4, 6, and 8 filter rounds on the CIFAR-100 to CIFAR-10 task are 92.18%, 94.92%, and 89.84%, respectively. Those results indicate that increasing the number of filtering rounds can enhance robustness against overwriting attacks to a certain extent. However, when the number of filtering rounds exceeds a certain threshold, the robustness may be slightly compromised due to the reduction in the number of parameters. Moreover, we observe the same trend as shown in Table V, when $\eta>=0.1$, the original watermark detection rate of NeuralMark falls below the security boundary, but the model performance is significantly impaired, demonstrating the attack is not effective.
Fine-tuning Attacks. Table X lists the results of fine-tuning attacks across distinct filtering rounds. We find that NeuralMark maintains a watermark detection rate of 100% across all cases, with negligible impact on the model performance.
TABLE X: Comparison of resistance to fine-tuning attacks with distinct filter rounds using ResNet-18. Watermark detection rates are omitted as they all reach 100%.
Fine-tuning Clean 4 Filters 6 Filters 8 Filters
CIFAR-100 to CIFAR-10 93.21 93.74 93.01 93.55
CIFAR-10 to CIFAR-100 72.17 71.67 72.68 72.27
@@
Pruning Attacks. Fig. 5 shows the results of pruning attacks on the CIFAR-10 and CIFAR-100 datasets using ResNet-18 across different filtering rounds. As can be seen, as the number of filtering rounds increases, the robustness of NeuralMark in resisting pruning attacks exhibits a slight decline. One reason is that increasing the number of filter rounds reduces the number of filtered parameters, leading to a smaller average pooling window size, which affects the robustness against pruning attacks.
@@
Figure 5: Comparison of resistance to pruning attacks with distinct filter rounds on CIFAR-10 (a) and CIFAR-100 (b) using ResNet-18 at various pruning ratios.
Question 8. How does the hashed watermark filter affect the overlap rate between the model owner and the adversary?
To analyze the effect of the hashed watermark filter on the overlap rate between the model owner and the adversary, we generate five counterfeit watermarks and calculate the overlap ratio between the parameters filtered by those and the original watermark. As shown in Fig. 6, the overlap rate decreases towards zero with more filtering rounds, indicating that watermark filtering enhances the confidentiality of the watermarked parameters.
VI-E Additional Analysis
Question 9. How does NeuralMark affect the parameter distribution?
To assess the influence of NeuralMark on the parameter distribution, Fig. 7 present the parameter distributions on the CIFAR-100 dataset with ResNet-18 and ViT-B/16 architectures. As can be seen, the parameter distributions of Clean and NeuralMark are nearly indistinguishable. Thus, it is challenging for adversaries to detect the embedded watermarks within the model. More parameter distribution results are provided Fig. 12 in Appendix C.
Figure 7: Comparison of parameter distributions on the CIFAR-100 dataset using distinct architectures: (a) ResNet-18, and (b) ViT-B/16.
TABLE XI: Comparison of the effects of average pooling on resistance to fine-tuning and pruning attacks using ResNet-18. Values (%) inside and outside the bracket are watermark detection rate and classification accuracy, respectively.
Method CIFAR-100 to CIFAR-10 Fine-tuning (Learning Rate) CIFAR-100 Pruning (Pruning Ratio)
0.001 0.005 0.01 40% 60% 80%
NeuralMark (w/o AP) 93.26 (100) 92.20 (100) 90.68 (81.64) 71.82 (90.62) 57.50 (78.51) 16.14 (69.92)
NeuralMark 93.74 (100) 92.25 (100) 91.25 (96.87) 69.86 (100) 43.88 (99.21) 9.85 (99.21)
TABLE XII: Comparison of average time cost (in seconds) on CIFAR-100 using ResNet-18. Here, $R$ is the number of filtering rounds.
Time (s) 23.60 24.49 24.94 25.01 25.19 24.34 47.43 35.17
Question 10. How does NeuralMark influence the performance convergence?
To examine the impact of NeuralMark on model performance convergence, Fig. 8 shows the results on the CIFAR-100 dataset with ResNet-18 and ViT-B/16 architectures. We find that the performance curves of Clean and NeuralMark exhibit a similar trend of change and are closely aligned, indicating that NeuralMark does not negatively affect the convergence of model performance. More performance convergence results are offered in Fig. 13 in Appendix D.
Figure 8: Comparison of performance convergence on the CIFAR-100 dataset using distinct architectures: (a) ResNet-18, and (b) ViT-B/16.
Question 11. How does average pooling impact NeuralMark?
To verify the efficacy of average pooling, we compare NeuralMark with its variant without average pooling, i.e, NeuralMark w/o AP. As shown in Table XI, both versions resist fine-tuning attacks at lower learning rates. However, at a learning rate of $0.01$, the detection rate for NeuralMark (w/o AP) drops to $81.64\%$, below the security boundary, while NeuralMark maintains at $96.87\%$. In addition, the detection rate of NeuralMark (w/o AP) rapidly declines with increasing pruning rates, reaching $69.92\%$ at an $80\%$ pruning rate, while NeuralMark achieves $99.21\%$. Those results confirm that average pooling enhances resistance to both fine-tuning and pruning attacks.
Question 12. Does NeuralMark impose a significant additional computational burden during training?
Table XII list the average time cost (in seconds) per training epoch over five epochs on the CIFAR-100 dataset using ResNet-18. NeuralMark’s running time is comparable to that of Clean and VanillaMark, highlighting the efficiency of NeuralMark. Also, NeuralMark outperforms GreedyMark in terms of speed due to GreedyMark’s reliance on costly sorting operations for parameter selection. Moreover, NeuralMark demonstrates significantly faster running times compared to VoteMark, as it avoids the multiple rounds of watermark embedding loss calculations required by VoteMark. Those results highlight the superior efficiency of NeuralMark.
Question 13. How do the watermark embedding layers impact the model’s performance?
To investigate the impact of watermark embedding layers on the model performance, we randomly choose four individual layers and all layers from ResNet-18 for watermark embedding. Table XIII presents the results on the CIFAR-100 dataset, showing that embedding different layers or all layers does not significantly affect the model performance.
TABLE XIII: Comparison of classification accuracy (%) on different watermarking layers on CIFAR-100 using ResNet-18. Here, Layers 1-4 denote randomly chosen layers, while All Layer refers to all layers. Watermark detection rates are omitted as they all reach 100%.
Watermarking Layer Layer 1 Layer 2 Layer 3 Layer 4 All Layer
Accuracy 76.51 76.68 76.30 76.73 75.86
Question 14. What is the effect of varying the watermark length on model performance?
To evaluate the influence of watermark length on the model performance, we set watermark lengths to 64, 128, 256, 512, 1024, and 2048, respectively. Table XIV lists the results on the CIFAR-100 dataset using ResNet-18, indicating that NeuralMark can achieve a $100\%$ detection rate with various watermark lengths while preserving nearly lossless model performance.
TABLE XIV: Comparison of classification accuracy (%) for distinct watermark lengths on CIFAR-100 using ResNet-18. Watermark detection rates are omitted as they all reach $100\%$.
Watermark Length 64 128 256 512 1024 2048
Accuracy 75.84 75.90 76.46 76.18 76.51 76.27
VII Conclusion
In this article, we present NeuralMark, a white-box method designed to protect model ownership. At the core of NeuralMark is a hashed watermark filter, which utilizes a hash function to generate an irreversible binary watermark from a secret key, subsequently employing this watermark as a filter to select model parameters for embedding. This simple design cleverly intertwines the embedding parameters with the hashed watermarks, providing robust protection against both forging and overwriting attacks. Moreover, the incorporation of average pooling provides resilience against fine-tuning and pruning attacks. We provide a theoretical analysis of NeuralMark’s security boundary. Extensive experiments on various datasets, architectures, and tasks confirm NeuralMark’s effectiveness and robustness. In the future, we plan to investigate how the proposed hashed watermark filter can be generalized and incorporated into a wider range of watermarking methods.
Appendix A Proof for Proposition V.1
Proposition V.1. Under the assumption that the hash function produces uniformly distributed outputs [38], for a model watermarked by NeuralMark with a watermark tuple $\{\mathbf{K},\mathbf{b}\}$, where $\mathbf{b}=\mathcal{H}(\mathbf{K})$, if an adversary attempts to forge a counterfeit watermark tuple $\{\mathbf{K}^{\prime},\mathbf{b}^{\prime}\}$ such that $\mathbf{b}^{\prime}=\mathcal{H}(\mathbf{K}^{\prime})$ and $\mathbf{K}^{\prime}\neq\mathbf{K}$, then the probability of achieving a watermark detection rate of at least $\rho$ (i.e., $\geq\rho$) is upper-bounded by $\frac{1}{2^{n}}\sum_{i=0}^{n-\lceil\rho n\rceil}\binom{n}{i}$.
$\begin{split}\Pr\big{[}X\geq\lceil\rho n\rceil\big{]}&=\sum_{i=\lceil\rho n\rceil}^{n}\binom{n}{i}\left(\frac{1}{2}\right)^{i}\left(\frac{1}{2}\right)^{n-i}\\ &=\frac{1}{2^{n}}\sum_{i=\lceil\rho n\rceil}^{n}\binom{n}{i}=\frac{1}{2^{n}}\sum_{i=0}^{n-\lceil\rho n\rceil}\binom{n}{i}.\end{split}$ (4)
Figure 9: Comparison of resistance to pruning attacks at various pruning ratios on CIFAR-100 using AlexNet and ResNet-18: (a) NeuralMark, (b) VanillaMark, (c) GreedyMark, and (d) VoteMark.
Figure 10: Comparison of resistance to pruning attacks at various pruning ratios on Caltech-101 using AlexNet and ResNet-18: (a) NeuralMark, (b) VanillaMark, (c) GreedyMark, and (d) VoteMark.
Figure 11: Comparison of resistance to pruning attacks at various pruning ratios on Caltech-256 using AlexNet and ResNet-18: (a) NeuralMark, (b) VanillaMark, (c) GreedyMark, and (d) VoteMark.
Appendix B Additional Results of Pruning Attacks
Fig. 10-11 provide the results from pruning attacks conducted on the CIFAR-100, Caltech-101, and Caltech-256 datasets, respectively. As can be seen, as the pruning ratio increases, the performance of NeuralMark degrades while the detection rate remains nearly 100%. This indicates NeuralMark’s robustness against pruning attacks. Those results collectively suggest NeuralMark exhibits superior robustness in resisting pruning attacks compared to other methods.
Appendix C Additional Results of Parameter Distribution
Fig. 12 provides additional parameter distributions for various architectures on the CIFAR-100 dataset. As can be seen, the parameter distributions of Clean and NeuralMark closely align in each architecture. Those results further demonstrate the secrecy of NeuralMark.
Appendix D Additional Results of Performance Convergence
Fig. 13 presents additional performance convergence plots for various architectures on the CIFAR-100 dataset. Across all architectures, the performance curves of Clean and NeuralMark exhibit similar trends and are closely aligned, further confirming that NeuralMark does not negatively affect performance convergence.
Figure 12: Comparison of parameter distributions on CIFAR-100 with distinct architectures: (a) AlexNet, (b) ResNet-18, (c) ResNet-34, (d) ViT-B/16, (e) VGG-16, (f) MobileNet-V3-L, (g) GoogLeNet, and (h) Swin-V2-B.
Figure 13: Comparison of model performance convergence on CIFAR-100 with distinct architectures: (a) AlexNet, (b) ResNet-18, (c) ResNet-34, (d) ViT-B/16, (e) VGG-16, (f) MobileNet-V3-L, (g) GoogLeNet, and (h) Swin-V2-B.

### v2

@@ F.1 Overwriting Attacks
F.2 Fine-tuning Attacks on All Model Parameters
We conduct a fine-tuning attack by updating all model parameters. The results are reported in Table 8. As shown, NeuralMark consistently demonstrates superior robustness, indicating that fine-tuning attacks cannot effectively remove watermarks embedded with NeuralMark. Notably, the model performance of all methods improves substantially under this setting. Specifically, on the CIFAR-10 to CIFAR-100 task using ResNet-18, NeuralMark achieves an accuracy of 71.67%, significantly higher than the 49.77% accuracy obtained when only the watermark embedding layers and classifier are fine-tuned (see Table 6). Similar trends are observed across other methods. Those results suggest that fine-tuning only the watermark-specific layers and classifier makes it difficult to preserve model performance.
Furthermore, to assess the impact of the learning rate $\eta$ in fine-tuning attacks, we evaluate settings of $\eta=0.01$ and $0.1$, in comparison to the default value of $0.001$. Those attacks are performed by updating all model parameters, allowing the model to retain effective model performance and thus posing a more practical threat. Table 9 presents the results on the CIFAR-100 to CIFAR-10 task using ResNet-18. At $\eta=0.01$, NeuralMark consistently achieves a higher watermark detection rate than other methods, demonstrating its robustness. At $\eta=0.1$, the model performance of all methods drops significantly, indicating that overly aggressive fine-tuning disrupts model functionality and renders the attack ineffective.
Table 8: Comparison of resistance to fine-tuning attacks against all layers. Values (%) inside and outside the bracket are watermark detection rate and classification accuracy, respectively.
@@ F.2 Fine-tuning Attacks on All Model Parameters
Table 9: Comparison of resistance to fine-tuning attacks against all layers at various learning rates ($\eta$) using ResNet-18. Values (%) inside and outside the bracket are the watermark detection rate and classification accuracy, respectively.
Fine-tuning $\eta$ Clean NeuralMark VanillaMark GreedyMark VoteMark
CIFAR-100 to CIFAR-10 0.001 93.21 93.74(100) 93.29(100) 93.21(100) 93.59(100)
0.01 91.22 90.41(97.65) 92.04(58.59) 90.91(82.81) 88.95(50.39)
0.1 69.05 78.71(58.20) 76.43(51.17) 82.34(62.10) 84.72(50.00)
F.3 Pruning Attacks
Figures 6-9 provide all the results from pruning attacks conducted on the CIFAR-10, CIFAR-100, Caltech-101, and Caltech-256 datasets, respectively. As can be seen, as the pruning ratio increases, the performance of NeuralMark degrades while the detection rate remains nearly 100%. This indicates NeuralMark’s robustness against pruning attacks. Those results collectively suggest NeuralMark exhibits superior robustness in resisting pruning attacks compared to other methods.
(a) NeuralMark
(b) VanillaMark
(c) GreedyMark
(d) VoteMark
@@ F.3 Pruning Attacks
Figure 6: Comparison of resistance to pruning attacks under various pruning ratios on CIFAR-10 using AlexNet and ResNet-18.
(a) NeuralMark
(b) VanillaMark
(c) GreedyMark
(d) VoteMark
Figure 7: Comparison of resistance to pruning attacks at various pruning ratios on CIFAR-100 using AlexNet and ResNet-18.
(a) NeuralMark
(b) VanillaMark
(c) GreedyMark
(d) VoteMark
Figure 8: Comparison of resistance to pruning attacks at various pruning ratios on Caltech-101 using AlexNet and ResNet-18.
(a) NeuralMark
(b) VanillaMark
(c) GreedyMark
(d) VoteMark
Figure 9: Comparison of resistance to pruning attacks at various pruning ratios on Caltech-256 using AlexNet and ResNet-18.
F.4 Parameter Distribution
Figure 10 provides additional parameter distributions for various architectures on the CIFAR-100 dataset. As can be seen, the parameter distributions of Clean and NeuralMark closely align in each architecture. Those results further demonstrate the secrecy of NeuralMark.
(a) AlexNet
(b) ResNet-18
(c) ResNet-34
(d) ViT-B/16
(e) VGG-16
(f) MobileNet-V3-L
(g) GoogLeNet
(h) Swin-V2-B
Figure 10: Comparison of parameter distributions on CIFAR-100 with distinct architectures.
F.5 Performance Convergence
Figure 11 presents additional performance convergence plots for various architectures on the CIFAR-100 dataset. Across all architectures, the performance curves of Clean and NeuralMark exhibit similar trends and are closely aligned, further confirming that NeuralMark does not negatively affect performance convergence.
(a) AlexNet
(b) ResNet-18
(c) ResNet-34
(d) ViT-B/16
(e) VGG-16
(f) MobileNet-V3-L
(g) GoogLeNet
(h) Swin-V2-B
Figure 11: Comparison of model performance convergence across distinct architectures on CIFAR-100.
F.6 Watermark Embedding Layers
To investigate the impact of watermark embedding layers on the model performance, we randomly choose four individual layers and all layers from ResNet-18 for watermark embedding. Table 10 presents the results on CIFAR-100, showing that embedding different layers or all layers does not significantly affect the model performance.
@@ F.6 Watermark Embedding Layers
Table 10: Comparison of classification accuracy (%) on different watermarking layers on CIFAR-100 using ResNet-18. Here, Layers 1-4 denote randomly chosen layers, while All Layers refers to all layers. Watermark detection rates are omitted as they all reach 100%.
Watermarking Layer Layer 1 Layer 2 Layer 3 Layer 4 All Layers
Accuracy 76.51 76.68 76.30 76.73 75.86
F.7 Watermark Length
To evaluate the influence of watermark length on the model performance, we set watermark lengths to 64, 128, 256, 512, 1024, and 2048, respectively. Table 11 lists the results on the CIFAR-100 dataset using ResNet-18, indicating that NeuralMark can achieve a $100\%$ detection rate with various watermark lengths while preserving nearly lossless model performance.
Table 11: Comparison of classification accuracy (%) for distinct watermark lengths on CIFAR-100 using ResNet-18. Watermark detection rates are omitted as they all reach $100\%$.
Watermark Length 64 128 256 512 1024 2048
Accuracy 75.84 75.90 76.46 76.18 76.51 76.27
F.8 Training Efficiency
Table 12 lists the average time cost (in seconds) per training epoch over five epochs on the CIFAR-100 dataset using ResNet-18. NeuralMark’s running time is comparable to that of Clean and VanillaMark, highlighting the efficiency of NeuralMark. Additionally, NeuralMark outperforms GreedyMark in terms of speed, as GreedyMark relies on costly sorting operations for parameter selection. Moreover, NeuralMark demonstrates significantly faster running times compared to VoteMark, as it avoids the multiple rounds of watermark embedding loss calculations required by VoteMark. Those results highlight the superior efficiency of NeuralMark.
Table 12: Comparison of average time cost (in seconds) on CIFAR-100 using ResNet-18. Here, $R$ denotes the number of filtering rounds.
Time (s) 23.60 24.49 24.94 25.01 25.19 24.34 47.43 35.17
Appendix G G. Further Analysis on Filtering Rounds
To assess the influence of the number of filtering rounds on NeuralMark’s effectiveness and robustness in resisting various attacks, we conduct additional experiments using 6 and 8 filters, compared to NeuralMark’s default setting of 4 filters. We omit forging attacks, since the hashed watermark filter is intrinsically resistant to such attacks.
G.1 Fidelity Evaluation
Table 13 presents the impact of watermark embedding on the model performance across distinct filtering rounds. The results demonstrate that NeuralMark, even with varying filtering rounds, has a minimal effect on the model performance while successfully embedding watermarks.
Table 13: Comparison of classification accuracy (%) with various distinct filter rounds on CIFAR-10 and CIFAR-100 using ResNet-18. Watermark detection rates are omitted as they all reach 100%.
@@ G.1 Fidelity Evaluation
G.2 Robustness Evaluation
Fine-tuning Attacks on All Model Parameters
Table 14 reports the results of fine-tuning attacks across distinct filtering rounds. The attacks are performed by updating all model parameters with a learning rate of 0.001. As shown, NeuralMark maintains a watermark detection rate of 100% across all filtering rounds, with negligible impact on model performance.
Table 14: Comparison of resistance to fine-tuning attacks with distinct filter rounds using ResNet-18. Watermark detection rates are omitted as they all reach 100%.
Fine-tuning Clean 4 Filters 6 Filters 8 Filters
CIFAR-100 to CIFAR-10 93.21 93.74 93.01 93.55
CIFAR-10 to CIFAR-100 72.17 71.67 72.68 72.27
Overwriting Attacks
Table 15 lists the results of overwriting attacks across distinct filtering rounds. From the results, we find that when the number of filtering rounds is set to 6, NeuralMark exhibits superior robustness compared to 4 and 8 filter rounds. Specifically, at $\eta=0.01$, the original watermark detection rates for 4, 6, and 8 filter rounds on the CIFAR-100 to CIFAR-10 task are 92.18%, 94.92%, and 89.84%, respectively. Those results indicate that increasing the number of filtering rounds can enhance robustness against overwriting attacks to a certain extent. However, when the number of filtering rounds exceeds a certain threshold, the robustness may be slightly compromised due to the reduction in the number of parameters.
Table 15: Comparison of resistance to overwriting attacks at various trade-off hyper-parameters ($\lambda$) and learning rates ($\eta$) with distinct filtering rounds using ResNet-18. Values (%) inside and outside the bracket are watermark detection rate and classification accuracy, respectively.
@@ G.2 Robustness Evaluation
Pruning Attacks
@@
Figure 12 shows the results of pruning attacks on CIFAR-10 and CIFAR-100 using ResNet-18 across different filtering rounds. As can be seen, as the number of filtering rounds increases, the robustness of NeuralMark in resisting pruning attacks exhibits a slight decline. One potential reason is that increasing the number of filter rounds reduces the number of filtered parameters, leading to a smaller average pooling window size, which affects the robustness against pruning attacks to some extent.
(a) CIFAR-10
(b) CIFAR-100
@@
Figure 12: Comparison of resistance to pruning attacks with distinct filter rounds on CIFAR-10 and CIFAR-100 using ResNet-18 at various pruning ratios.